A,B,C can complete a work in 2 hours . if A does the job alone in 6 hours and B in 5 hours , how long will it take for C to do the job alone
Answer by richwmiller(17219)

2/6+2/5+2/c=1
c = 15/2=7.5 hours to do the job alone
a does 1/3 or 5/15
b does 6/15
and c must do 4/15 of the job
c does 2/15 an hour